# PruningBench: A Comprehensive Benchmark of Structural Pruning

### v3

@@ 4 Benchmark Results and Discussions
2x FPGM [20] N/A 1 78.588 69.248 -9.34 10.365 M (47.01%) 0.937s N/A
Random [22] N/A 2 78.588 68.810 -9.778 9.305 M (42.20%) 0.888s N/A
LAMP [23] N/A 3 78.588 68.724 -9.864 10.169 M (46.12%) 1.284s N/A
MagnitudeL1 [21] N/A 4 78.588 68.602 -9.986 10.375 M (47.05%) 1.005s N/A
MagnitudeL2 [21] N/A 5 78.588 68.316 -10.272 10.346 M (46.92%) 0.995s N/A
OBD-Hessian [15] N/A 6 78.588 67.514 -11.074 10.334 M (46.87%) 6m40s N/A
Taylor [19] N/A 7 78.588 67.400 -11.188 10.468 M (47.47%) 27.634s N/A
CP [24] N/A 7 78.588 67.400 -11.188 10.334 M (46.87%) 15m4s N/A
ThiNet [25] N/A 8 78.588 63.914 -14.674 6.439 M (29.20%) 3m17s N/A
MagnitudeL2 [21] GrowingReg [11] 1 78.588 68.715 -9.873 10.359 M (46.98%) 0.995s 5h10m31s
MagnitudeL2 [21] GroupNorm [15] 2 78.588 68.594 -9.994 10.363 M (47.00%) 0.995s 5h21m21s
MagnitudeL2 [21] GroupLASSO [61] 3 78.588 68.350 -10.238 10.360 M (46.98%) 0.995s 5h15m13s
3x MagnitudeL1 [21] N/A 1 78.588 63.120 -15.468 6.57 M (29.79%) 1.005s N/A
LAMP [23] N/A 2 78.588 62.538 -16.050 6.08 M (27.57%) 1.284s N/A
MagnitudeL2 [21] N/A 3 78.588 62.342 -16.246 6.37 M (28.89%) 0.995s N/A
Taylor [19] N/A 4 78.588 61.582 -17.006 6.62 M (30.01%) 27.634s N/A
FPGM [20] N/A 5 78.588 60.660 -17.928 5.701 M (25.85%) 0.937s N/A
CP [24] N/A 6 78.588 56.626 -21.962 6.778 M (30.74%) 15m4s N/A
OBD-Hessian [15] N/A 7 78.588 54.796 -23.792 6.39 M (28.98%) 6m40s N/A
ThiNet [25] N/A 8 78.588 49.654 -28.934 5.113 M (23.19%) 3m17s N/A
Random [22] N/A 9 78.588 44.654 -33.954 4.95 M (22.45%) 0.888s N/A
MagnitudeL2 [21] GrowingReg [11] 1 78.588 62.608 -15.980 6.57 M (29.81%) 0.995s 5h10m31s
MagnitudeL2 [21] GroupNorm [15] 2 78.588 61.716 -16.872 6.88 M (31.20%) 0.995s 5h21m21s
MagnitudeL2 [21] GroupLASSO [61] 3 78.588 61.340 -17.248 6.57 M (29.13%) 0.995s 5h15m13s
4x MagnitudeL1 [21] N/A 1 78.588 59.950 -18.638 5.06 M (22.93%) 1.005s N/A
MagnitudeL2 [21] N/A 2 78.588 59.082 -19.506 4.89 M (22.15%) 0.995s N/A
Taylor [19] N/A 3 78.588 57.650 -20.938 4.80 M (21.76%) 27.634s N/A
LAMP [23] N/A 4 78.588 55.750 -22.838 4.32 M (19.57%) 1.284s N/A
FPGM [20] N/A 5 78.588 48.258 -30.33 3.25 M (14.74%) 0.937 N/A
OBD-Hessian [15] N/A 6 78.588 36.600 -41.988 4.25 M (19.27%) 6m40s N/A
CP [24] N/A 7 78.588 42.574 -36.014 5.253 M (23.82%) 15m4s N/A
ThiNet [25] N/A 8 78.588 28.422 -50.166 2.669 M (12.10%) 3m17s N/A
Random [22] N/A 9 78.588 27.722 -50.866 2.76 M (12.54%) 0.888s N/A
MagnitudeL2 [21] GrowingReg [11] 1 78.588 59.630 -18.958 4.56 M (20.66%) 0.995s 5h10m31s
MagnitudeL2 [21] GroupLASSO [61] 2 78.588 57.312 -21.276 4.59 M (20.81%) 0.995s 5h15m13s
MagnitudeL2 [21] GroupNorm [15] 3 78.588 56.446 -22.142 4.77 M (21.62%) 0.995s 5h21m21s
PruningBench now has completed $645$ model pruning experiments (i.e., getting $645$ pruned models), yielding $13$ leaderboards (9 on CIFAR, 3 on ImageNet, and one on COCO). For space considerations, here we present the leaderboard results of ResNet50 on CIFAR100 in Table 2, and the leaderboard results of ViT-small on ImageNet in Table 3. For more leaderboards, please refer to Table A3 to A15 in SUPP or our benchmark website at http://pruning.vipazoo.cn.
@@ 4.1 Benchmark Results
Overall Results. In general, no single methods consistently outperform the others across all settings and tasks. Nonetheless, weight norm-based methods, such as MagnitudeL1 and MagnitudeL2, typically exhibit superior performance and yield more reliable results, ranking within the top $5$ in most rankings while maintaining computational efficiency. This is followed by BNScale, FPGM, Taylor, and OBD-C, which also show commendable results in various scenarios. Other methodologies may not exhibit significant overall advantages, but may perform well in specific situations. Now we provide more detailed analyses of the leaderboard results by answering the following questions.
Q1: What is the impact of the model architectures on the leaderboard rankings?
@@
Q2: What is the impact of the speedup ratio on the leaderboard rankings?
@@
Observation: Different methods can exhibit varying rankings under different speedup ratios. MagnitudeL1, MagnitudeL2, BNScale, and LAMP slightly improve in ranking as the speedup ratio increases, indicating a certain level of pruning resilience. Conversely, FPGM, ThiNet, and Hrank tend to experience a decline in rankings as the speedup ratio increases. OBD-Hessian and CP methods show relatively stable performance, with minimal ranking shifts across speedup ratios. Taylor and OBD-C, however, display more erratic behavior, with their rankings sometimes significantly rising or falling depending on the architecture and speedup ratios.
@@
Q3: Which methods are more efficient in terms of computation time?
Observation: Obviously, sparsifying-stage methods are significantly more computation expensive than pruning-stage methods due to the cumbersome sparse learning process. In general, in our experiments sparsifying-stage methods takes about $1.33\sim 2$ times longer time than pruning-stage methods. For pruning-stage methods, data-driven importance criteria [13, 24, 18, 9, 19, 25, 22, 62, 15], particularly those involving non-parallel operations [9, 15, 24, 25], consume longer pruning time compared to data-free methods. For example, OBD-Hessian [15] computes gradients separately for each sample. Thinet [25] and HRank [9] determine the importance of each output channel of each layer individually. These techniques are well-suited for one-shot pruning, where importance scores are calculated only once, followed by pruning the network to achieve the target pruning ratio. However, when applied to iterative pruning, the computation time increases significantly as the importance scores need to be calculated every iteration.
Q4: How do sparsity regularizers improve the performance of prunned model?
Observation: Sparsity regularizers [11, 15, 26, 61] aim to induce sparsity in network parameters, rendering redundant parameters proximate to zero or outright zero, thereby facilitating pruning based on importance criteria. However, in practical applications, we find that sparsity regularizers do not necessarily improve performance and only show significant effects in specific scenarios. Notably, across all experiments utilizing sparsity regularizers (refer to Table A3-A15 in the SUPP), only $57.30\%$ showcase positive performance improvements with sparsity regularization. Among these techniques, BNScale delivers the most favorable outcomes, with a $77.78\%$ probability of enhancing performance, followed by GroupLASSO with a $65.38\%$ likelihood of improvement. Conversely, GroupNorm and GrowingReg demonstrate lower effectiveness overall, with improvement probabilities of $42.31\%$ and $45.83\%$, respectively. Nonetheless, these methods excel in particular architectural settings. GrowingReg, for instance, excels in the ViT architecture, manifesting notable performance enhancements across all speedup ratios, while other techniques improve ViT performance less than half of the time. GroupNorm, on the other hand, is better suited for VGG models, exhibiting a $66.67\%$ probability of performance enhancement, a significant improvement compared to its performance in other architectures. A drawback of sparsity regularizers is the necessity for meticulous tuning tailored to each scenario, i.e., the optimal hyperparameters vary across diverse model-dataset configurations. Please refer to the SUPP for more details of the hyperparameters.
Q5: How consistent are the CIFAR rankings with the ImageNet rankings?
Observation: In comparison to pruning CIFAR100-trained models, pruning ImageNet-trained models (of the same architecture as CIFAR100-trained ones) typically results in more accuracy deterioration. Meanwhile, ImageNet-trained models are more sensitive to the speedup ratios. However, for the same model architecture, the leaderboard rankings on CIFAR are highly consistent with that on ImageNet (see Table A5, A8, A13-14 in SUPP for details). For example, when pruning ResNet models, MagnitudeL1, MagnitudeL2, BNScale, and Taylor methods consistently rank within the top five on both CIFAR100 and ImageNet, whereas LAMP and Hrank consistently rank low on the list. These observations indicate that pruning methods showcase a degree of consistency across datasets with the same model. In situations where computational resources are constrained, the utilization of smaller datasets for assessing pruning methodologies, followed by the application of the top-ranked techniques to prune larger models, emerges as a viable approach.
4.2 More Discussions
Local pruning, global pruning versus protected global pruning. Table 4 presents the results of ResNet50 pruned by MagnitudeL2 on ImageNet, with the three pruning schemes. Results on more models are provided in SUPP. From these results, we make the following two conclusions. (1) At low speedup ratios, global pruning and the proposed protected global pruning exhibit comparable performance, both surpassing local pruning. It can be attributed to the assumption in local pruning, which assigns equal importance to each group and applies the same pruning ratio to each group, overlooking group differences. To address this, some previous works such as [63, 21, 25, 13] propose sensitivity analysis in order to estimate the pruning ratio that should be applied to particular layers [19]. (2) In contrast, at higher speedup ratios, protected global pruning outperforms both local pruning and global pruning. An examination of the network architecture after global pruning uncover instances of layer collapse, where nearly all channels of a network layer are eliminated, rendering the network untrainable and severely impairing performance.
@@ 4.2 More Discussions
Figure 2: The parameter curves of different models pruned by different importance criteria on CIFAR100 dataset. Details can be referred to in Tables A5, A8, and A11 in the SUPP.
@@
Parameters versus FLOPS. Some prior works [10, 23, 41, 42, 43, 44, 45, 28, 29] employ the number of parameters as the performance metric of pruned models. Here we discuss the correlation between parameters and the computation cost, i.e., FLOPS. As evidenced in Table 2 and 3, different methods may yield significantly different numbers of pruned parameters at the identical speedup ratios, indicating unequal contributions of various parameters to the computational overhead. Specifically, as depicted in Figure 2, at the same speedup ratio, models with fewer total parameters have more parameters in their initial blocks. Here we provide a brief theoretical insight on this phenomenon. For a convolutional layer $\mathbf{W}\in\mathbb{R}^{\hat{N}NK^{2}}$, the input tensor $\mathbf{I}\in\mathbb{R}^{NHW}$, and the output tensor $\mathbf{O}\in\mathbb{R}^{\hat{N}\hat{H}\hat{W}}$, the computational complexity of this layer can be denoted by $\mathbf{I}$ is $O(N\hat{N}\hat{H}\hat{W}{K}^{2})$. The average computational contribution of each parameter is thus $O(\hat{H}\hat{W})$, which implies a positive linear correlation between the computation overhead and the feature map resolution. As CNN models usually progressively scale down the spatial resolution of feature maps as layers deepen, the method that prioritizes the reduction of shallow parameters can effectively decrease computational costs while minimizing the parameter count. However, the same method can exhibit varying preferences when pruning different architectures. For example, OBD-Hessian removes numerous parameters when pruning ResNet18 and ResNet50 (see Table 2 and Table A3 to Table A8 in SUPP), indicating a preference for pruning later layers. However, when pruning VGG19, it removes much fewer parameters, suggesting a focus on earlier layers (see Table A9 to Table A11 in SUPP).
@@
CNNs versus ViTs. CNNs and ViTs present diverse characteristics and demonstrate distinct behavious in structural pruning. For instance, in the case of CNN architectures, owing to the aforementioned relationship between parameters and computational overhead, there can be large differences in the number of parameters pruned by different methods at the identical speedup ratios (see Table 2 and other CNN experiments in SUPP). However, for the ViT-small experiments in Table 3, the differences in the number of parameters among different methods at the same speedup ratio are small. This phenomenon arises from the fixed-shaped flattened tensors that characterize the output feature maps of ViT, ensuring a consistent contribution of parameters to computational overhead across distinct layers. Therefore, in contrast to CNN, using the number of parameters as the performance metric for pruning ViTs can also lead to reliable conclusions, thanks to the nearly linear correlation between parameters and computational cost.
@@
Another crucial aspect of Vision Transformers (ViTs) is the intricate interconnection of the patch embedding layer with other layers. The output dimension of the patch embedding layer plays a pivotal role in determining the input dimension for all attention layers, making ViT pruning particularly sensitive to this layer. Additionally, ViT necessitates pruning same dimensions for different attention heads, thereby increasing the implementation complexity. Moreover, in comparison to ResNet50, ViT-small, with a similar model size, suffer from more accuracy loss. As depicted in Table 5, at the same speedup ratio, the accuracy loss of ViT-small is several times greater than that of ResNet50. The experimental result aligns with the general consensus in prior literature [64, 65, 66, 67, 68] that ViT models are harder to compress while preserving accuracy compared to their classic convolutional counterparts. For further comparisons, please consult Table 3 and Table A14 in the SUPP.
@@
Method applicability. The applicability of different pruning methods exhibits considerable variance. Most methods are tailored for CNNs, presenting obstacles when adapting them to alternative architectural designs. For instance, HRank [9] determines channel importance based on the rank of feature map corresponding to each channel, which is incompatible with architectures like ViTs. Since ViTs output flattened tensors, pruning through this method is unfeasible. Analogous challenges emerge with Batch Normalization (BN)-based techniques [26, 37, 38, 30, 39], which rely on batch normalization layers for importance score. Consequently, these methods can not be directly applied to architectures without batch normalization layers. In contrast to the previously mentioned approaches, techniques based on weight normalization [21, 69, 23] and weight similarity [20, 70, 48, 32, 33] exhibit minimal constraints and can be seamlessly integrated into diverse architectural frameworks.

### v5

@@ 5 Benchmark Results and Discussions
2x FPGM N/A 1 78.588 69.248 -9.34 10.365 M (47.01%) 0.937s N/A
Random N/A 2 78.588 68.810 -9.778 9.305 M (42.20%) 0.888s N/A
LAMP N/A 3 78.588 68.724 -9.864 10.169 M (46.12%) 1.284s N/A
MagnitudeL1 N/A 4 78.588 68.602 -9.986 10.375 M (47.05%) 1.005s N/A
MagnitudeL2 N/A 5 78.588 68.316 -10.272 10.346 M (46.92%) 0.995s N/A
OBD-Hessian N/A 6 78.588 67.514 -11.074 10.334 M (46.87%) 6m40s N/A
Taylor N/A 7 78.588 67.400 -11.188 10.468 M (47.47%) 27.634s N/A
CP N/A 7 78.588 67.400 -11.188 10.334 M (46.87%) 15m4s N/A
ThiNet N/A 8 78.588 63.914 -14.674 6.439 M (29.20%) 3m17s N/A
MagnitudeL2 GrowingReg 1 78.588 68.715 -9.873 10.359 M (46.98%) 0.995s 5h10m31s
MagnitudeL2 GroupNorm 2 78.588 68.594 -9.994 10.363 M (47.00%) 0.995s 5h21m21s
MagnitudeL2 GroupLASSO 3 78.588 68.350 -10.238 10.360 M (46.98%) 0.995s 5h15m13s
3x MagnitudeL1 N/A 1 78.588 63.120 -15.468 6.57 M (29.79%) 1.005s N/A
LAMP N/A 2 78.588 62.538 -16.050 6.08 M (27.57%) 1.284s N/A
MagnitudeL2 N/A 3 78.588 62.342 -16.246 6.37 M (28.89%) 0.995s N/A
Taylor N/A 4 78.588 61.582 -17.006 6.62 M (30.01%) 27.634s N/A
FPGM N/A 5 78.588 60.660 -17.928 5.701 M (25.85%) 0.937s N/A
CP N/A 6 78.588 56.626 -21.962 6.778 M (30.74%) 15m4s N/A
OBD-Hessian N/A 7 78.588 54.796 -23.792 6.39 M (28.98%) 6m40s N/A
ThiNet N/A 8 78.588 49.654 -28.934 5.113 M (23.19%) 3m17s N/A
Random N/A 9 78.588 44.654 -33.954 4.95 M (22.45%) 0.888s N/A
MagnitudeL2 GrowingReg 1 78.588 62.608 -15.980 6.57 M (29.81%) 0.995s 5h10m31s
MagnitudeL2 GroupNorm 2 78.588 61.716 -16.872 6.88 M (31.20%) 0.995s 5h21m21s
MagnitudeL2 GroupLASSO 3 78.588 61.340 -17.248 6.57 M (29.13%) 0.995s 5h15m13s
4x MagnitudeL1 N/A 1 78.588 59.950 -18.638 5.06 M (22.93%) 1.005s N/A
MagnitudeL2 N/A 2 78.588 59.082 -19.506 4.89 M (22.15%) 0.995s N/A
Taylor N/A 3 78.588 57.650 -20.938 4.80 M (21.76%) 27.634s N/A
LAMP N/A 4 78.588 55.750 -22.838 4.32 M (19.57%) 1.284s N/A
FPGM N/A 5 78.588 48.258 -30.33 3.25 M (14.74%) 0.937 N/A
OBD-Hessian N/A 6 78.588 36.600 -41.988 4.25 M (19.27%) 6m40s N/A
CP N/A 7 78.588 42.574 -36.014 5.253 M (23.82%) 15m4s N/A
ThiNet N/A 8 78.588 28.422 -50.166 2.669 M (12.10%) 3m17s N/A
Random N/A 9 78.588 27.722 -50.866 2.76 M (12.54%) 0.888s N/A
MagnitudeL2 GrowingReg 1 78.588 59.630 -18.958 4.56 M (20.66%) 0.995s 5h10m31s
MagnitudeL2 GroupLASSO 2 78.588 57.312 -21.276 4.59 M (20.81%) 0.995s 5h15m13s
MagnitudeL2 GroupNorm 3 78.588 56.446 -22.142 4.77 M (21.62%) 0.995s 5h21m21s
PruningBench now has completed $645$ model pruning experiments (i.e., getting $645$ pruned models), yielding $13$ leaderboards (9 on CIFAR, 3 on ImageNet, and one on COCO). For space considerations, here we present the leaderboard results of ResNet50 on CIFAR100 in Table 2, and the leaderboard results of ViT-small on ImageNet in Table 3. For more leaderboards, please refer to our online pruning platform http://pruning.vipazoo.cn.
@@ 5.1 Benchmark Results
Overall Results. In general, no single method consistently outperforms the others across all settings and tasks. Nonetheless, weight norm-based methods, such as MagnitudeL1 and MagnitudeL2, typically exhibit superior performance and yield more reliable results, ranking within the top $5$ in most rankings while maintaining computational efficiency. This is followed by BNScale, FPGM, Taylor, and OBD-C, which also show commendable results in various scenarios. Other methodologies may not exhibit significant overall advantages, but may perform well in specific situations. Now we provide more detailed analyses of the leaderboard results by answering the following questions.
Q1: How do the model architectures impact leaderboard rankings?
@@
Q2: What is the impact of the speedup ratio on leaderboard rankings?
@@
Observation: Different methods can exhibit varying rankings under different speedup ratios. MagnitudeL1, MagnitudeL2, BNScale, and LAMP slightly improve in ranking as the speedup ratio increases, indicating a certain level of pruning resilience. Conversely, FPGM, ThiNet, and Hrank tend to experience a decline in rankings as the speedup ratio increases. OBD-Hessian and CP methods show relatively stable performance, with minimal ranking shifts across speedup ratios. Taylor and OBD-C, however, display more erratic behavior, with their rankings sometimes rising or falling significantly depending on the architecture and speedup ratios.
Q3: Which methods are more efficient in terms of computation time?
Observation: Obviously, sparsifying-stage methods are significantly more computation expensive than pruning-stage methods due to the cumbersome sparse learning process. In general, in our experiments sparsifying-stage methods take about $1.33\sim 2$ times longer time than pruning-stage methods. For pruning-stage methods, data-driven importance criteria [14, 25, 19, 10, 20, 26, 23, 62, 16], particularly those involving non-parallel operations [10, 16, 25, 26], consume longer pruning time compared to data-free methods. For example, OBD-Hessian [16] computes gradients separately for each sample. Thinet [26] and HRank [10] determine the importance of each output channel of each layer individually. These techniques are well-suited for one-shot pruning, where importance scores are calculated only once, followed by pruning the network to achieve the target pruning ratio. However, when applied to iterative pruning, the computation time increases significantly as importance scores need to be calculated every iteration.
Q4: How do regularizers improve the performance of prunned model?
Observation: Sparsity regularizers [12, 16, 27, 55] aim to induce sparsity in network parameters, rendering redundant parameters proximate to zero or outright zero, thereby facilitating pruning based on importance criteria. However, in practical applications, we find that sparsity regularizers do not necessarily improve performance and only show significant effects in specific scenarios. Notably, across all experiments utilizing sparsity regularizers, only $57.30\%$ showcase positive performance improvements with sparsity regularization. Among these techniques, BNScale delivers the most favorable outcomes, having a $77.78\%$ probability of enhancing performance, followed by GroupLASSO with a $65.38\%$ likelihood of improvement. Conversely, GroupNorm and GrowingReg demonstrate lower effectiveness overall, with improvement probabilities of $42.31\%$ and $45.83\%$, respectively. Nonetheless, these methods excel in particular architectural settings. GrowingReg, for instance, excels in the ViT architecture, manifesting notable performance enhancements across all speedup ratios, while other techniques improve ViT performance less than half of the time. GroupNorm, on the other hand, is better suited for VGG models, exhibiting a $66.67\%$ probability of performance enhancement, a significant improvement compared to its performance in other architectures. A drawback of sparsity regularizers is the necessity for meticulous tuning tailored to each scenario, i.e., the optimal hyperparameters vary across diverse model-dataset configurations. Please refer to the A.1.4 for more details of the hyperparameters.
Q5: How consistent are the CIFAR rankings with ImageNet rankings?
Observation: In comparison to pruning CIFAR100-trained models, pruning ImageNet-trained models (of the same architecture as CIFAR100-trained ones) typically results in greater accuracy deterioration. Meanwhile, ImageNet-trained models are more sensitive to the speedup ratios. However, for the same model architecture, the leaderboard rankings on CIFAR are highly consistent with those on ImageNet (see leaderboards in https://github.com/HollyLee2000/PruningBench for details). For example, when pruning ResNet models, MagnitudeL1, MagnitudeL2, BNScale, and Taylor methods consistently rank within the top five on both CIFAR100 and ImageNet, whereas LAMP and Hrank consistently rank low on the list. These observations indicate that pruning methods showcase a degree of consistency across datasets with the same model. In situations where computational resources are constrained, the utilization of smaller datasets for assessing pruning methodologies, followed by the application of the top-ranked techniques to prune larger models, emerges as a viable approach.
5.2 More Discussions
Local pruning, global pruning versus protected global pruning. Table 4 presents the results of ResNet50 pruned by MagnitudeL2 on ImageNet, with the three pruning schemes. Results on more models are provided in A.4. From these results, we get the following two conclusions. (1) At low speedup ratios, global pruning and the proposed protected global pruning exhibit comparable performance, both surpassing local pruning. It can be attributed to the assumption in local pruning, which assigns equal importance to each group and applies the same pruning ratio to each group, overlooking group differences. To address this, some previous works such as [71, 22, 26, 14] propose sensitivity analysis in order to estimate the pruning ratio that should be applied to particular layers [20]. (2) In contrast, at higher speedup ratios, protected global pruning outperforms both local pruning and global pruning. An examination of the network architecture after global pruning uncover instances of layer collapse, where nearly all channels of a network layer are eliminated, rendering the network untrainable and severely impairing performance.
@@ 5.2 More Discussions
Figure 2: The parameter curves of different models pruned by different importance criteria on CIFAR100 dataset. Details can be referred to the github leaderboards.
@@
Parameters versus FLOPS. Some prior works [11, 24, 42, 43, 44, 45, 59, 29, 30] employ the number of parameters as the performance metric of pruned models. Here we discuss the correlation between parameters and the computation cost, i.e., FLOPS. As evidenced in Table 2 and 3, different methods may yield significantly different numbers of pruned parameters at the identical speedup ratios, indicating unequal contributions of various parameters to the computational overhead. Specifically, as depicted in Figure 2, at the same speedup ratio, models with fewer total parameters have more parameters in their initial blocks. Here we provide a brief theoretical insight on this phenomenon. For a convolutional layer $\mathbf{W}\in\mathbb{R}^{\hat{N}NK^{2}}$, the input tensor $\mathbf{I}\in\mathbb{R}^{NHW}$, and the output tensor $\mathbf{O}\in\mathbb{R}^{\hat{N}\hat{H}\hat{W}}$, the computational complexity of this layer can be denoted by $\mathbf{I}$ is $O(N\hat{N}\hat{H}\hat{W}{K}^{2})$. The average computational contribution of each parameter is thus $O(\hat{H}\hat{W})$, which implies a positive linear correlation between the computation overhead and the feature map resolution. As CNN models usually progressively scale down the spatial resolution of feature maps as layers deepen, the method that prioritizes the reduction of shallow parameters can effectively decrease computational costs while minimizing the parameter count. However, the same method can exhibit varying preferences when pruning different architectures. For example, OBD-Hessian removes numerous parameters when pruning ResNet18 and ResNet50 (see Table 2), indicating a preference for pruning later layers. However, when pruning VGG19, it removes much fewer parameters, suggesting a focus on earlier layers.
@@
CNNs versus ViTs. CNNs and ViTs present diverse characteristics and demonstrate distinct behavious in structural pruning. For instance, in the case of CNN architectures, owing to the aforementioned relationship between parameters and computational overhead, there can be large differences in the number of parameters pruned by different methods at the identical speedup ratios (see Table 2 and other CNN experiments in A.4). However, for the ViT-small experiments in Table 3, the differences in the number of parameters among different methods at the same speedup ratio are small. This phenomenon arises from the fixed-shaped flattened tensors that characterize the output feature maps of ViT, ensuring a consistent contribution of parameters to computational overhead across distinct layers. Therefore, in contrast to CNN, using the number of parameters as the performance metric for pruning ViTs can also lead to reliable conclusions, thanks to the nearly linear correlation between parameters and computational cost.
@@
Another crucial aspect of Vision Transformers (ViTs) is the intricate interconnection of the patch embedding layer with other layers. The output dimension of the patch embedding layer plays a pivotal role in determining the input dimension for all attention layers, making ViT pruning particularly sensitive to this layer. Additionally, ViT necessitates pruning same dimensions for different attention heads, thereby increasing the implementation complexity. Moreover, in comparison to ResNet50, ViT-small with a similar model size, suffers from more accuracy loss. As depicted in Table 5, at the same speedup ratio, the accuracy loss of ViT-small is several times greater than that of ResNet50. The experimental result aligns with the general consensus in prior literature [72, 73, 74, 75, 76] that ViT models are harder to compress while preserving accuracy compared to their classic convolutional counterparts. For further comparisons, please consult Table 3.
@@
Method applicability. The applicability of different pruning methods exhibits considerable variance. Most methods are tailored for CNNs, presenting obstacles when adapting them to alternative architectural designs. For instance, HRank [10] determines channel importance based on the rank of feature map corresponding to each channel, which is incompatible with architectures like ViTs. Since ViTs output flattened tensors, pruning through this method is unfeasible. Analogous challenges emerge with Batch Normalization (BN)-based techniques [27, 38, 39, 31, 40], which rely on batch normalization layers for importance score. Consequently, these methods can not be directly applied to architectures without batch normalization layers. In contrast to the previously mentioned approaches, techniques based on weight normalization [22, 56, 24] and weight similarity [33, 21, 34, 57] exhibit minimal constraints and can be seamlessly integrated into diverse architectural frameworks.